40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2589     Entry time: Tue Sep 15 15:13:44 2020
Author: Anchal 
Type: Notes 
Category: Environment 
Subject: HEPA filters switched on. Measurement stopped. 

I have switched on HEPA filters to high, both on top of the main table and on top of the flow bench.


Continuous measurement is stopped hereby. This experiment is finished.

ELOG V3.1.3-